# Weyl Law and convergence in the classical limit for min-max nonlocal minimal surfaces

### v3

@@ 5.4.1 Estimates for $s$-minimal surfaces with bounded index
The second term is the main one. To obtain a bound for $\Big{|}\textnormal{Per}(E_{i};\mathcal{C}_{j})-(1-s_{i})\textnormal{Per}_{s_{i}}(E_{i};\mathcal{C}_{j})\Big{|}$, we can argue as in Proposition 5.26, using $F_{j}^{-1}:C_{j}\subset\mathbb{R}^{n}\to\mathcal{C}_{j}$ as a chart parametrisation and the fact that $F_{j}$ is a $(1+\delta)$-biLipschitz equivalence. The only difference is that $C_{j}$ is not a cylinder, so that the $F_{j}(E_{i})$ cannot be exactly written as a union of subgraphs inside $C_{j}$; on the other hand, the $F_{j}$ are the restriction of geodesic normal coordinate maps $\widetilde{F}_{j}:B_{r_{j}}\to\mathcal{B}_{r_{j}}$ to the $C_{j}$, so that (up to choosing $r_{j}$ small enough) the $\widetilde{F}_{j}(E_{i})$ are indeed a union of subgraphs $A_{k}$ inside $\mathcal{B}_{r_{j}}$. We can then decompose $F_{j}(E_{i})=\cup_{k}(A_{k}\cap C_{j})$, where (as argued at the beginning) the $A_{k}\cap C_{j}$ are quantitatively transversal the $\partial C_{j}$. This allows us to apply [16, Lemma 11] to each of the $A_{k}\cap C_{j}$, exactly as in the proof of Proposition 5.26. Anyhow, we conclude that
@@
By Theorem 5.31, we have the convergence of a subsequence of the $\partial E_{i}$ to a limit stationary integral varifold $V$ in the varifold sense. Moreover, by Theorem 5.35, this varifold is supported on a smooth surface $\Sigma$. Let $\Sigma_{1},...,\Sigma_{m}$ denote the connected components of $\Sigma$; then, by the Constancy Theorem (see [48, Theorem 41.1]), the $\Sigma_{k}$ are minimal surfaces and there are integers $n_{k}$ such that $V=\sum_{k=1}^{m}n_{k}[\Sigma_{k}]$, where $\Sigma_{k}]$ denotes the integer rectifiable varifold associated to the smooth surface $\Sigma_{k}$. Combining this with Proposition 5.34, we deduce that
@@
We start with the following simple lemma, an analogue of [42, Lemma 1] and which shows that the $\mathfrak{p}$-widths are continuous (more precisely, locally Lipschitz) with respect to the metric:
@@
The proof is just as in [38, Proposition 3.1], which uses the minimal surfaces associated to the Almgren-Pitts spectrum and the Weyl Law associated to them. One simply need substitute the continuity of the $\mathfrak{p}$-widths in [38, Lemma 2.1] by Lemma 6.1, the attainability of the min-max widths in [38, Proposition 2.2] by Corollary 1.5, and the Weyl Law from [40] by Theorem 1.3. ∎
@@
The proof in [42], which again uses the minimal surfaces and Weyl Law associated to the Almgren-Pitts spectrum, can be replicated in our case. The result in [42, Lemma 1] on the Lipschitz continuity of the $\mathfrak{p}$-widths is replaced by Lemma 6.1 from the present paper, the attainability of the min-max widths in [38, Proposition 2.2] (which is used in the proof of [42, Lemma 2]) is replaced by Corollary 1.5, and the Weyl Law from [40] is replaced by Theorem 1.3. ∎
@@
The existence of a critical point $u_{\varepsilon,\mathfrak{p}}$ with energy $c_{\varepsilon,s}$ follows from applying a min-max theorem to the functional $\mathcal{E}_{\varepsilon,s}$ with $\widetilde{\mathcal{F}}_{\mathfrak{p}}$ as the set of min-max families, exactly as in the proof of [18, Theorem 3.4]. The fact that $u_{\varepsilon,\mathfrak{p}}$ has index bounded by $\mathfrak{p}$ comes from the fact that min-max critical points of families of the type (35) satisfy this upper index bound; this is thanks to the Hausdorff measure condition $\mathcal{H}^{\mathfrak{p}}(A)<\infty$, as described in [39], which intuitively corresponds to considering sets $A$ with at most $\mathfrak{p}$ parameters in the min-max.
Finally, the upper and lower energy bounds in (37) follow the same strategy as in the proof of [18, Theorem 3.2]. To be precise, in [18] we considered the family
@@
On the other hand, the upper bounds were shown by constructing an explicit $f:\mathbb{S}^{\mathfrak{p}}\to H^{s/2}(M)\setminus{0}$ such that all elements in $A=f(\mathbb{S}^{\mathfrak{p}})\in\mathcal{G}_{\mathfrak{p}}$ have low energy. If the Hausdorff measure condition $\mathcal{H}^{\mathfrak{p}}(A)<\infty$ holds, then $A\in\mathcal{F}_{\mathfrak{p}}$ and thus the same set $A$ works in our case. Unfortunately, the only reasonable way of showing the measure condition is to see that $f:\mathbb{S}^{\mathfrak{p}}\to H^{s/2}(M)\setminus{0}$ is Lipschitz, which is not true for the function $f$ considered in [18, Theorem 3.2] – in fact, it is only $\beta$-Holder continuous for an intermediate exponent $\beta\in(0,1)$ in general. We now explain the construction and how to adapt it to our case to obtain a Lipschitz map.
@@
which after taking square roots on both sides proves that $a\mapsto u_{a,\delta}$ is a Lipschitz map from $\mathbb{S}^{\mathfrak{p}}$ to $H^{s/2}(\mathcal{B}_{r_{\mathfrak{p}}})$. This concludes the proof in the case $M=\mathcal{B}_{r_{\mathfrak{p}}}$. As mentioned at the beginning, in general one actually decomposes $M$ as a union of $\mathfrak{p}$ subsets which (passing to normal coordinates) are of this type; see the proof of [18, Theorem 3.2]. The additional interactions between the respective subsets can simply be bounded directly by the fractional perimeters of the subsets, including in the proof of the Lipschitz property above in the general case. ∎
@@ Appendix C Some Riemannian computations
As usual, $g_{ij}$ will denote the coefficients of the metric $g$ in the coordinates given by $\varphi$. We want to embed isometrically a piece of our manifold $M$ inside a torus with perturbed metric (from its original flat one). For that, let $\eta:\mathbb{R}^{n}\to\mathbb{R}$ be a standard cutoff, with $\chi_{\mathcal{B}_{1}}\leq\eta\leq\chi_{\mathcal{B}_{2}}$. Consider the $n$-dimensional flat torus $(\mathbb{T}^{n},\delta_{ij})$ of side length $8$, obtained from the hypercube of side length $8$ centered at $0\in\mathbb{R}^{n}$ by identifying opposite faces, and let $\varphi^{\prime}:\mathbb{R}^{n}\to\mathbb{T}^{n}$ denote the quotient map defining $\mathbb{T}^{n}$. Define a new Riemannian manifold $M^{\prime}=(\mathbb{T}^{n},g^{\prime})$, where $g^{\prime}_{ij}:=\eta g_{ij}+(1-\eta)\delta_{ij}$ is just the usual flat metric except on $\varphi^{\prime}(\mathcal{B}_{2})\subset\mathbb{T}^{n}$. It is obvious then that $M^{\prime}$ satisfies the flatness assumptions ${\rm FA}_{\ell}(M^{\prime},g^{\prime},1,p_{0}^{\prime}=0,\varphi^{\prime})$, since by construction $g_{ij}\equiv g^{\prime}_{ij}$ in $\mathcal{B}_{1}$ in the coordinates given by $\varphi$ and $\varphi^{\prime}$. Thanks to this, as recorded in [17, Lemma 2.17] the heat kernels of both manifolds in $\mathcal{B}_{1/2}$ are actually quantitatively close, in the sense that
@@
For $h^{\prime}=h$, i.e. without quantifying the result on the metric, this is a standard result obtained by constructing the heat kernel on a closed manifold using some parametrix approach. The fact that the coefficients in the expansion can be quantified in terms of the metric, in the sense that the result holds uniformly for $h^{\prime}$ satisfying the assumptions in the present proposition, follows from a careful inspection of the version of the approach mentioned above which uses normal coordinates to construct the parametrix. This is a standard way of constructing the parametrix, which the reader can find for example in [45, Chapter 3]. In this approach, one first constructs locally a parametrix working on normal coordinates around every point of the manifold, by explicitly solving certain ODEs in polar coordinates to define an approximate heat kernel (parametrix) of the form
@@
Finally, one constructs the heat kernel on $M\times M$ iteratively, essentially by considering the error $LH^{(k)}$ of the approximate heat kernel $H^{(k)}$ at a certain step $k$ to vanishing under the heat operator $L:=\partial_{t}-\Delta$, convolving this error with the approximate heat kernel $H^{(k)}$ itself, and adding this quantity to $H^{(k)}$ to obtain a better approximation $H^{(k+1)}$; see [45, page 98] for the formal computation. This procedure converges to the true (unique) solution $H_{N,h^{\prime}}$, and the difference/error between $H^{(0)}$ and $H_{N,h^{\prime}}$ is then given by an explicit infinite sum made of iterated convolutions between $H^{(0)}$ and $LH^{(0)}$. This error difference (together with its first $\lfloor l-n/2\rfloor\geq 2$ spatial derivatives) is then shown to be uniformly bounded by $Ct^{l+1}$, see [45, Lemma 3.18], for all $t\in(0,1]$, which gives our desired result. Here $C$ is a constant which depends on $H^{(0)}$, $LH^{(0)}$ and the metric tensor of $(N,h^{\prime})$, which shows that $C$ is uniformly controlled for all $h^{\prime}$: since $H^{(0)}$ is explicitly constructed in terms of the metric in normal coordinates (as explained at the beginning), both $H^{(0)}$ and $LH^{(0)}$ (and, of course, the metric tensor itself $h^{\prime}$ as well) are uniformly controlled for all $h^{\prime}$ satisfying our assumptions. ∎
@@
Finally, $\nabla H_{t}$ and $\nabla E_{t}$ are uniformly bounded for large times: it is clear for $E_{t}$ from its explicit formula, and a constant bound for the derivatives of $H_{t}$ is obtained in the proof of [17, Theorem 2.13], by combining Lemma D.2 with [17, Proposition 2.19]. Therefore,

### v4

@@ 5.4.1 Estimates for $s$-minimal surfaces with bounded index
The second term is the main one. To obtain a bound for $\Big|\textnormal{Per}(E_{i};\mathcal{C}_{j})-(1-s_{i})\textnormal{Per}_{s_{i}}(E_{i};\mathcal{C}_{j})\Big|$, we can argue as in Proposition 5.26, using $F_{j}^{-1}:C_{j}\subset\mathbb{R}^{n}\to\mathcal{C}_{j}$ as a chart parametrisation and the fact that $F_{j}$ is a $(1+\delta)$-biLipschitz equivalence. The only difference is that $C_{j}$ is not a cylinder, so that the $F_{j}(E_{i})$ cannot be exactly written as a union of subgraphs inside $C_{j}$; on the other hand, the $F_{j}$ are the restriction of geodesic normal coordinate maps $\widetilde{F}_{j}:B_{r_{j}}\to\mathcal{B}_{r_{j}}$ to the $C_{j}$, so that (up to choosing $r_{j}$ small enough) the $\widetilde{F}_{j}(E_{i})$ are indeed a union of subgraphs $A_{k}$ inside $\mathcal{B}_{r_{j}}$. We can then decompose $F_{j}(E_{i})=\cup_{k}(A_{k}\cap C_{j})$, where (as argued at the beginning) the $A_{k}\cap C_{j}$ are quantitatively transversal the $\partial C_{j}$. This allows us to apply [17, Lemma 11] to each of the $A_{k}\cap C_{j}$, exactly as in the proof of Proposition 5.26. Anyhow, we conclude that
@@
By Theorem 5.31, we have the convergence of a subsequence of the $\partial E_{i}$ to a limit stationary integral varifold $V$ in the varifold sense. Moreover, by Theorem 5.35, this varifold is supported on a smooth surface $\Sigma$. Let $\Sigma_{1},...,\Sigma_{m}$ denote the connected components of $\Sigma$; then, by the Constancy Theorem (see [51, Theorem 41.1]), the $\Sigma_{k}$ are minimal surfaces and there are integers $n_{k}$ such that $V=\sum_{k=1}^{m}n_{k}[\Sigma_{k}]$, where $\Sigma_{k}]$ denotes the integer rectifiable varifold associated to the smooth surface $\Sigma_{k}$. Combining this with Proposition 5.34, we deduce that
@@
We start with the following simple lemma, an analogue of [44, Lemma 1] and which shows that the $\mathfrak{p}$-widths are continuous (more precisely, locally Lipschitz) with respect to the metric:
@@
The proof is just as in [40, Proposition 3.1], which uses the minimal surfaces associated to the Almgren-Pitts spectrum and the Weyl Law associated to them. One simply need substitute the continuity of the $\mathfrak{p}$-widths in [40, Lemma 2.1] by Lemma 6.1, the attainability of the min-max widths in [40, Proposition 2.2] by Corollary 1.5, and the Weyl Law from [42] by Theorem 1.3. ∎
@@
The proof in [44], which again uses the minimal surfaces and Weyl Law associated to the Almgren-Pitts spectrum, can be replicated in our case. The result in [44, Lemma 1] on the Lipschitz continuity of the $\mathfrak{p}$-widths is replaced by Lemma 6.1 from the present paper, the attainability of the min-max widths in [40, Proposition 2.2] (which is used in the proof of [44, Lemma 2]) is replaced by Corollary 1.5, and the Weyl Law from [42] is replaced by Theorem 1.3. ∎
@@
The existence of a critical point $u_{\varepsilon,\mathfrak{p}}$ with energy $c_{\varepsilon,s}$ follows from applying a min-max theorem to the functional $\mathcal{E}_{\varepsilon,s}$ with $\widetilde{\mathcal{F}}_{\mathfrak{p}}$ as the set of min-max families, exactly as in the proof of [19, Theorem 3.4]. The fact that $u_{\varepsilon,\mathfrak{p}}$ has index bounded by $\mathfrak{p}$ comes from the fact that min-max critical points of families of the type (35) satisfy this upper index bound; this is thanks to the Hausdorff measure condition $\mathcal{H}^{\mathfrak{p}}(A)<\infty$, as described in [41], which intuitively corresponds to considering sets $A$ with at most $\mathfrak{p}$ parameters in the min-max.
Finally, the upper and lower energy bounds in (37) follow the same strategy as in the proof of [19, Theorem 3.2]. To be precise, in [19] we considered the family
@@
On the other hand, the upper bounds were shown by constructing an explicit $f:\mathbb{S}^{\mathfrak{p}}\to H^{s/2}(M)\setminus{0}$ such that all elements in $A=f(\mathbb{S}^{\mathfrak{p}})\in\mathcal{G}_{\mathfrak{p}}$ have low energy. If the Hausdorff measure condition $\mathcal{H}^{\mathfrak{p}}(A)<\infty$ holds, then $A\in\mathcal{F}_{\mathfrak{p}}$ and thus the same set $A$ works in our case. Unfortunately, the only reasonable way of showing the measure condition is to see that $f:\mathbb{S}^{\mathfrak{p}}\to H^{s/2}(M)\setminus{0}$ is Lipschitz, which is not true for the function $f$ considered in [19, Theorem 3.2] – in fact, it is only $\beta$-Holder continuous for an intermediate exponent $\beta\in(0,1)$ in general. We now explain the construction and how to adapt it to our case to obtain a Lipschitz map.
@@
which after taking square roots on both sides proves that $a\mapsto u_{a,\delta}$ is a Lipschitz map from $\mathbb{S}^{\mathfrak{p}}$ to $H^{s/2}(\mathcal{B}_{r_{\mathfrak{p}}})$. This concludes the proof in the case $M=\mathcal{B}_{r_{\mathfrak{p}}}$. As mentioned at the beginning, in general one actually decomposes $M$ as a union of $\mathfrak{p}$ subsets which (passing to normal coordinates) are of this type; see the proof of [19, Theorem 3.2]. The additional interactions between the respective subsets can simply be bounded directly by the fractional perimeters of the subsets, including in the proof of the Lipschitz property above in the general case. ∎
@@ Appendix C Some Riemannian computations
As usual, $g_{ij}$ will denote the coefficients of the metric $g$ in the coordinates given by $\varphi$. We want to embed isometrically a piece of our manifold $M$ inside a torus with perturbed metric (from its original flat one). For that, let $\eta:\mathbb{R}^{n}\to\mathbb{R}$ be a standard cutoff, with $\chi_{\mathcal{B}_{1}}\leq\eta\leq\chi_{\mathcal{B}_{2}}$. Consider the $n$-dimensional flat torus $(\mathbb{T}^{n},\delta_{ij})$ of side length $8$, obtained from the hypercube of side length $8$ centered at $0\in\mathbb{R}^{n}$ by identifying opposite faces, and let $\varphi^{\prime}:\mathbb{R}^{n}\to\mathbb{T}^{n}$ denote the quotient map defining $\mathbb{T}^{n}$. Define a new Riemannian manifold $M^{\prime}=(\mathbb{T}^{n},g^{\prime})$, where $g^{\prime}_{ij}:=\eta g_{ij}+(1-\eta)\delta_{ij}$ is just the usual flat metric except on $\varphi^{\prime}(\mathcal{B}_{2})\subset\mathbb{T}^{n}$. It is obvious then that $M^{\prime}$ satisfies the flatness assumptions ${\rm FA}_{\ell}(M^{\prime},g^{\prime},1,p_{0}^{\prime}=0,\varphi^{\prime})$, since by construction $g_{ij}\equiv g^{\prime}_{ij}$ in $\mathcal{B}_{1}$ in the coordinates given by $\varphi$ and $\varphi^{\prime}$. Thanks to this, as recorded in [18, Lemma 2.17] the heat kernels of both manifolds in $\mathcal{B}_{1/2}$ are actually quantitatively close, in the sense that
@@
For $h^{\prime}=h$, i.e. without quantifying the result on the metric, this is a standard result obtained by constructing the heat kernel on a closed manifold using some parametrix approach. The fact that the coefficients in the expansion can be quantified in terms of the metric, in the sense that the result holds uniformly for $h^{\prime}$ satisfying the assumptions in the present proposition, follows from a careful inspection of the version of the approach mentioned above which uses normal coordinates to construct the parametrix. This is a standard way of constructing the parametrix, which the reader can find for example in [48, Chapter 3]. In this approach, one first constructs locally a parametrix working on normal coordinates around every point of the manifold, by explicitly solving certain ODEs in polar coordinates to define an approximate heat kernel (parametrix) of the form
@@
Finally, one constructs the heat kernel on $M\times M$ iteratively, essentially by considering the error $LH^{(k)}$ of the approximate heat kernel $H^{(k)}$ at a certain step $k$ to vanishing under the heat operator $L:=\partial_{t}-\Delta$, convolving this error with the approximate heat kernel $H^{(k)}$ itself, and adding this quantity to $H^{(k)}$ to obtain a better approximation $H^{(k+1)}$; see [48, page 98] for the formal computation. This procedure converges to the true (unique) solution $H_{N,h^{\prime}}$, and the difference/error between $H^{(0)}$ and $H_{N,h^{\prime}}$ is then given by an explicit infinite sum made of iterated convolutions between $H^{(0)}$ and $LH^{(0)}$. This error difference (together with its first $\lfloor l-n/2\rfloor\geq 2$ spatial derivatives) is then shown to be uniformly bounded by $Ct^{l+1}$, see [48, Lemma 3.18], for all $t\in(0,1]$, which gives our desired result. Here $C$ is a constant which depends on $H^{(0)}$, $LH^{(0)}$ and the metric tensor of $(N,h^{\prime})$, which shows that $C$ is uniformly controlled for all $h^{\prime}$: since $H^{(0)}$ is explicitly constructed in terms of the metric in normal coordinates (as explained at the beginning), both $H^{(0)}$ and $LH^{(0)}$ (and, of course, the metric tensor itself $h^{\prime}$ as well) are uniformly controlled for all $h^{\prime}$ satisfying our assumptions. ∎
@@
Finally, $\nabla H_{t}$ and $\nabla E_{t}$ are uniformly bounded for large times: it is clear for $E_{t}$ from its explicit formula, and a constant bound for the derivatives of $H_{t}$ is obtained in the proof of [18, Theorem 2.13], by combining Lemma D.2 with [18, Proposition 2.19]. Therefore,